# Representations of knot groups in $\textup{AGL}_{1}(\mathbb{C})$ and Alexander invariants

### v1

@@ Abstract.
This paper reinterprets Alexander-type invariants of knots via representation varieties of knot groups into the group $\textup{AGL}_{1}(\mathbb{C})$ of affine transformations of the complex line. In particular, we prove that the coordinate ring of the $\textup{AGL}_{1}(\mathbb{C})$-representation variety is isomorphic to the symmetric algebra of the Alexander module. This yields a natural interpretation of the Alexander polynomial as the singular locus of a coherent sheaf over $\mathbb{C}^{*}$, whose fibres correspond to quandle representation varieties of the knot quandle. As a by-product, we construct Topological Quantum Field Theories that provide effective computational methods and recover the Burau representations of braids. This theory offers a new geometric perspective on classical Alexander invariants and their functorial quantization.
@@ 1. Introduction
Since its inception in 1928, as introduced in [1], the Alexander polynomial has been one of the main invariants in knot theory. Over the past century, it has been extensively studied from several perspectives, playing a central role in connecting topological, algebraic and geometric approaches to the classification of knots. Alexander originally obtained it from a Dehn presentation of the knot group using a planar projection: a system of linear equations naturally arises, leading to a determinant that defines the polynomial. It was soon recognized that the Alexander polynomial is an invariant associated with the first homology group of the infinite cyclic cover of the knot complement. Consequently, it can be obtained from the Seifert matrix or via Fox calculus, as noted in reference texts on knot theory [9, 28, 32]. Conway later introduced the potential function [11] –now known as the Alexander-Conway polynomial– and demonstrated how to compute it recursively using skein relations, emphasizing the importance of local relations that were already present in Alexander’s original work. Subsequence advances in knot theory have further explored Alexander invariants from other points of view, such as the Vassiliev invariants [4, 45], the Reidemeister torsion [34, 43] or knot Floer homology [36], among many others.
The Alexander polynomial was long believed to be the only polynomial knot invariant until the discovery of the Jones and HOMFLY-PT polynomials [17, 24]. Years later, Witten’s seminal work [48] revolutionized knot theory, establishing Topological Quantum Field Theories (TQFTs) and various categorifications of algebraic knot invariants, notably Khovanov homology [29] for the Jones polynomial and its variation for the Alexander polynomial [39]. From the perspective of representation theory, a fruitful development has been the construction of Reshetikhin-Turaev invariants [38], which produce invariants of knots using representations of quantum groups. The Kauffman bracket and the Jones polynomial can thus be recovered from representations of certain quantum groups, such as $U_{q}(\mathfrak{sl}_{2})$, also leading to coloured variations.
@@
usually known as the knot group. In that case, very subtle information can be extracted from character varieties, such as in the case $G=\operatorname{SL}_{2}(\mathbb{C})$, where the character variety of the knot group gauges the hyperbolic structures on the knot complement [12, 13]. In this context, the theory of skein modules, which are related to quantum groups and quantum field theories, has been an intensive object of research in recent years [37]. Furthermore, if $\Gamma$ is the fundamental group of a surface, Turaev [44] showed that the Kauffman bracket skein algebra provides a quantization of the $\operatorname{SL}_{2}(\mathbb{C})$-character variety. Moreover, representations of this algebra generalize the Kauffman bracket invariant for links [5, 6]. For 3-manifolds, it has recently been proved that the Kauffman skein module is finite dimensional [19] and has been computed in certain cases [15]. We also note that in all these theories, the skein relation plays a fundamental role, both theoretically and computationally.
@@
It is worth pointing out that this idea of attaching a coherent sheaf to the ${\textup{AGL}_{1}}(\mathbb{C})$-representation variety has previously been explored by Hironaka in [23] from a cohomological point of view. In that work, equations defining the jumping loci for the first cohomology group of one-dimensional representations of a finitely presented group were computed via Fox calculus. These variations in dimension naturally induce a stratification of the character variety of a knot group, recovering the Alexander stratification in terms of coherent sheaves. Hironaka applied these results to establish a criterion for proving that certain finitely presented groups are not Kähler.
@@
[…]
@@
The functor ${\mathcal{A}}:{\mathcal{T}}\to\operatorname{Span}(\textbf{{SVectB}})$ must be understood as a Topological Quantum Field Theory (TQFT for short) from the category ${\mathcal{T}}$ of tangles with values in the category $\operatorname{Span}(\textbf{{SVectB}})$, in the sense of Atiyah [3] and Segal [40]. These TQFTs are the analogues of a functorial quantization of a given field theory, in which manifolds are quantized into a (Hilbert) vector space and bordisms are translated into a ‘path integral’ operation between the boundary Hilbert spaces. In the context of knot theory, it is customary to replace the target category of vector spaces by another monoidal category ($\operatorname{Span}(\textbf{{SVectB}})$ in our case) better adapted to the geometric problem but mimicking all the properties of a quantization procedure. These TQFT-based methods have been extensively studied in the literature due to their ability to capture and provide effective methods of computation of subtle invariants, such as (coloured) Jones or HOMFLY-PT polynomials, see [42] for a survey on the state-of-art. However, to the best of our knowledge no such TQFT has been previously constructed to quantize the Alexander polynomial in a functorial way. It is also worth noticing that the TQFT that we define ${\mathcal{A}}:{\mathcal{T}}\to\operatorname{Span}(\textbf{{SVectB}})$, produces a quantization of the Alexander polynomial in a different sense from the ‘categorification’ approaches, such as Khovanov-like methods [29, 39], which capture polynomial invariants in a distinct way, namely, by constructing a homology theory whose (graded) Euler characteristic recovers the polynomial.
@@
We finish this section by pointing out that the classical results re-proven in this work are just a few examples of the flexibility and simplicity that the developed TQFT-based framework confers. Thanks to the functor $\mathcal{A}_{t}:{\mathcal{T}}\to\operatorname{Span}(\operatorname{\mathbf{Vect}}_{{\mathbb{C}}})$, we can provide alternative proofs of well-known results in knot theory involving Alexander polynomials with a direct approach. Furthermore, we expect to exploit these techniques to produce new explicit descriptions of Alexander polynomial for certain families of knots, as will be presented in a separate note [18].
@@ 2. ${\textup{AGL}_{1}}$-representations of knots
[…]
@@
For simplicity, we will abbreviate ${\textup{AGL}_{1}}={\textup{AGL}_{1}}(\mathbb{C})$. Notice that algebraically, ${\textup{AGL}_{1}}={\mathbb{C}}^{*}\ltimes{\mathbb{C}}$, for the action of ${\mathbb{C}}^{*}$ on ${\mathbb{C}}$ by $t\cdot a=ta$.
@@
Analogously, each relation of type (II), $A_{j}^{-1}A_{i}A_{j}=A_{i+1}$, yields an analogous equation $-a_{j}t^{-1}+a_{j}+t^{-1}a_{i}=a_{i+1}$, equivalent to
@@
Regarding non-abelian affine representations of ${\Gamma}_{K}$, the following fundamental theorem originally proven by de Rham and Burde in [12, Section 6.1] plays a fundamental role. We restate it in terms of the fibration given in (9), and we will re-prove it in Section 3.1.
@@ 3.1. Fox calculus for the Alexander module
Throughout this section, let us fix a knot $K\subset S^{3}$ with complement $K^{c}=S^{3}-K$. The commutator subgroup $[\pi_{1}(K^{c}),\pi_{1}(K^{c})]\triangleleft\pi_{1}(K^{c})$ defines an infinite cyclic covering $K^{c}_{\infty}\rightarrow K^{c}$, with deck group isomorphic to $H_{1}(K^{c})=\pi_{1}(K^{c})/[\pi_{1}(K^{c}),\pi_{1}(K^{c})]\cong\mathbb{Z}=\langle t\rangle$, which is usually called the infinite cyclic covering of $K^{c}$. The cyclic action of $t$ endows each homology group $H_{k}(K^{c}_{\infty})$ with a $\mathbb{Z}[t,t^{-1}]$-module structure, giving rise to a collection of modules that are invariants of the knot, the Alexander modules of $K$.
@@
$E_{K}^{k}=\begin{cases}0,&\text{if }k\leq 0,\\ \text{ideal generated by the $(n-k)\times(n-k)$-minors of $A_{K}$},&\text{if }0<k\leq n-1,\\ {\mathbb{Z}}[t,t^{-1}],&\text{if }k\geq n.\end{cases}$
@@
The $k$-th Alexander polynomial $\Delta_{K}^{k}(t)$ of the knot $K$ is defined as the greatest common divisor of the elements of the ideal $E_{K}^{k}$, defined up to units ±$\pm t^{k}$, since ${\mathbb{Z}}[t,t^{-1}]$ is a unique factorization ring. In other words, $\Delta_{K}^{k}(t)$ is the generator of the smallest principal ideal containing $E_{K}^{k}$. Clearly, the first elementary ideal $E_{K}^{1}$ is a principal ideal, generated by a single polynomial $\Delta_{K}^{1}(t)$, also denoted by $\Delta_{K}(t)$, the Alexander polynomial, that appeared in Theorem 2.2.
@@
[28, Chapter 11] that $H_{1}(K^{c}_{\infty})$ is the ${\mathbb{Z}}[t,t^{-1}]$-module generated by $[x_{i}],i=1,\ldots,n$, under the abelianized Wirtinger relations:
@@
according to the Wirtinger presentation of the knot. As it is also stated in [28, section 11.2], the linear relations (13), (14) in ${\mathbb{Z}}[t,t^{-1}]$ of the Alexander module are precisely the ones arising in any affine representation, from (7) and (8).
@@ 3.2. ${\textup{AGL}_{1}}$-representation varieties and the Alexander module
$[{\mathcal{R}}_{{\textup{AGL}_{1}}}(K)]={\mathbb{L}}\left({\mathbb{L}}-1-\sum_{k=1}^{n-1}\left|S_{K}^{k}\right|\right)+\sum_{k=1}^{n-1}\left|S_{K}^{k}\right|{\mathbb{L}}^{k+1}={\mathbb{L}}({\mathbb{L}}-1)+\sum_{k=1}^{n-1}\left|S_{K}^{k}\right|({\mathbb{L}}^{k+1}-{\mathbb{L}}).$
@@ 3.3. The Alexander fibration as a singular vector bundle
[…]
and such that the changes of charts are fibrewise linear. A morphism between singular vector bundles $\pi:E\to X$ and $\pi^{\prime}:E^{\prime}\to X^{\prime}$ is a pair $(f,\tilde{f})$ of regular morphisms $f:X\to X^{\prime}$ and $\tilde{f}:E\to E^{\prime}$ such that the following diagram commutes
@@
Furthermore, as shown in [20, Proposition II.5.1], every coherent sheaf is uniquely determined by the finitely generated module $M_{K}=\mathcal{M}_{K}({\mathbb{C}}^{*})$ of global sections over the ring $\mathcal{O}_{{\mathbb{C}}^{*}}({\mathbb{C}}^{*})={\mathbb{C}}[t,t^{-1}]$. In fact, since $M_{K}$ is isomorphic to the ring of regular functions on ${\mathcal{R}}_{{\textup{AGL}_{1}}}(K)$, by Theorem 3.3 we get the following result.
@@
The coherent sheaf $\mathcal{M}_{K}$ has been previously studied by Hironaka in [23] in a different context, with an ad hoc construction by means of $1$-cocycles. In this direction, the Alexander fibration (15) coincides with the one studied in [23].

### v2

# Representations of knot groups in $\textrm{AGL}_{1}(\mathbb{C})$ and Alexander invariants

@@ Abstract.
This paper reinterprets Alexander-type invariants of knots via representation varieties of knot groups into the group $\textrm{AGL}_{1}(\mathbb{C})$ of affine transformations of the complex line. In particular, we prove that the coordinate ring of the $\textrm{AGL}_{1}(\mathbb{C})$-representation variety is isomorphic to the symmetric algebra of the Alexander module. This yields a natural interpretation of the Alexander polynomial as the singular locus of a coherent sheaf over $\mathbb{C}^{*}$, whose fibres correspond to quandle representation varieties of the knot quandle. As a by-product, we construct Topological Quantum Field Theories that provide effective computational methods and recover the Burau representations of braids. This theory offers a new geometric perspective on classical Alexander invariants and their functorial quantization.
@@ 1. Introduction
Since its inception in 1928, as introduced in [1], the Alexander polynomial has been one of the main invariants in knot theory. Over the past century, it has been extensively studied from several perspectives, playing a central role in connecting topological, algebraic and geometric approaches to the classification of knots. Alexander originally derived it from a Dehn presentation of the knot group using a planar projection: a system of linear equations naturally arises, leading to a determinant that defines the invariant. The Alexander polynomial was long believed to be the only polynomial knot invariant until the discovery of the Jones and HOMFLY-PT polynomials [23, 31]. Years later, Witten’s seminal work [60] revolutionized knot theory, establishing the framework for Topological Quantum Field Theories (TQFTs) and various categorifications of algebraic knot invariants, notably Khovanov homology [36] for the Jones polynomial and its variation for the Alexander polynomial [50]. From the perspective of representation theory, a fruitful development has been the construction of Reshetikhin-Turaev invariants [49], which produce invariants of knots using representations of quantum groups. The Kauffman bracket and the Jones polynomial can thus be recovered from representations of certain quantum groups, such as $U_{q}(\mathfrak{sl}_{2})$, leading also to coloured variations and several other insights.
@@
usually known as the knot group. In that case, very subtle information can be extracted from character varieties, such as in the case $G=\operatorname{SL}_{2}(\mathbb{C})$, where the character variety of the knot group gauges the hyperbolic structures on the knot complement [17, 18].
@@
${0}$${{\mathcal{R}}_{{\textup{AGL}_{1}}(\mathbb{C})}(\Gamma_{K})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0,}$
@@
The functor ${\mathcal{A}}:{\mathcal{T}}\to\operatorname{Span}(\textbf{{SVectB}})$ must be understood as a Topological Quantum Field Theory (TQFT for short) from the category $\mathcal{T}$} of tangles with values in the category $\operatorname{Span}(\textbf{{SVectB}})$, in the sense of Atiyah [3] and Segal [51]. These TQFTs are the analogues of a functorial quantization of a given field theory, in which manifolds are quantized into a (Hilbert) vector space and bordisms are translated into a ‘path integral’ operation between the boundary Hilbert spaces. In the context of knot theory, it is customary to replace the target category of vector spaces by another monoidal category ($\operatorname{Span}(\textbf{{SVectB}})$ in our case) better adapted to the geometric problem but mimicking all the properties of a quantization procedure. These TQFT-based methods have been extensively studied in the literature due to their ability to capture and provide effective methods of computation of subtle invariants, such as (coloured) Jones or HOMFLY-PT polynomials, see [53] for a survey on the state-of-art.
@@
We finish this section by pointing out that the classical results re-proven in this work are just a few examples of the flexibility and simplicity that the developed TQFT-based framework confers. Thanks to the functor $\mathcal{A}_{t}:{\mathcal{T}}\to\operatorname{Span}(\operatorname{\mathbf{Vect}}_{{\mathbb{C}}})$, we can provide alternative proofs of well-known results in knot theory involving Alexander polynomials with a direct approach. Furthermore, we expect to exploit these techniques to produce new explicit descriptions of the Alexander module for certain families of knots, which will be presented in a separate note [43].
Related work
The Alexander theory of a knot is a widely studied topic that has been analyzed through different viewpoints. It was soon recognized that the Alexander polynomial is an invariant associated with the first homology group of the infinite cyclic cover of the knot complement. Consequently, it can be obtained from the Seifert matrix or via Fox calculus, as noted in reference texts on knot theory [11, 35, 40]. Conway later introduced the potential function [16] –now known as the Alexander-Conway polynomial– and demonstrated how to compute it recursively using skein relations, emphasizing the importance of local relations that were already present in Alexander’s original work. Subsequence advances in knot theory have further explored Alexander invariants from other points of view, such as the Vassiliev invariants [4, 56], the Reidemeister torsion [45, 54] or knot Floer homology [47], among many others.
The relation of the Alexander polynomial with representation theory has also been explored in the literature. Back in the 60’s, de Rham [19] already reported that abelian representations of the $\operatorname{SL}_{2}(\mathbb{C})$-character variety appear precisely over the roots of the Alexander polynomial. This character variety actually plays a central role in the construction of other celebrated knot invariats, such as the A-polynomial [17] and the brand-new slope invariant [5]. This relation has been clarified through the theory of skein modules [48], by means of its relation with quantum groups and quantum field theories. Using the representation theory of surface groups, Turaev [55] showed that the Kauffman bracket skein algebra provides a quantization of the $\operatorname{SL}_{2}(\mathbb{C})$-character variety. Moreover, representations of this algebra generalize the Kauffman bracket invariant for links [7, 8]. For 3-manifolds, it has recently been proved that the Kauffman skein module is finite dimensional [26] and has been computed in certain cases [20].
However, to the best of our knowledge, it has not been reported in the literature such a clear relation between the geometry of the ${\textup{AGL}_{1}}(\mathbb{C})$-representation variety and Alexander modules as the one provided by our Theorem A. Perhaps the closest work in this direction was conducted by Hironaka [30], who exploited the idea of attaching a coherent sheaf from a purely cohomological point of view with the aim of identifying non-Kähler finitely generated groups. In that work, equations defining the jumping loci for the first cohomology group of one-dimensional representations of a finitely presented group were computed via Fox calculus. These variations in dimension naturally induce a stratification of the character variety of a knot group, recovering the Alexander stratification in terms of coherent sheaves. However, to the best of our understanding, this is a purely ad-hoc construction with no relation with the ${\textup{AGL}_{1}}(\mathbb{C})$-representation variety.
It is worth noticing that there exists in the literature several approaches to construct functorial quantum field theories for Alexander-type invariants. Apart from the groundbreaking work by Witten quantizing the Jones polynomial by means of Chern-Simons theory, perhaps the first construction of a TQFT for Alexander invariants was developed by Frohman and Nicas [24]. In this work, the authors constructed a functor $Z:\mathcal{C}\to\mathbb{P}\textup{{Ab}}$ from a suitable subcategory $\mathcal{C}$ of the category of $3$-dimensional bordisms to the category $\mathbb{P}\textup{{Ab}}$ of “projective” abelian groups, i.e. the usual category of abelian groups but with morphisms defined only up to re-scaling. This functor assigns to a connected surface $\Sigma$ the homology $H_{\star}({\mathcal{R}}_{\textup{U}(1)}(\pi_{1}(\Sigma)))$ of the associated $\textup{U}(1)$-representation variety, and is designed in such a way that it quantizes the intersection form. In this manner, this TQFT captures the Seifert matrix of a Seifert surface for a knot and thus the construction recovers the Alexander polynomial of the knot from the vacuum state associated to the bordism resulting from applying longitudinal surgery to the knot. It is worth mentioning that the use of $\textup{U}(1)$-representations is very natural in this context since $R_{\textup{U}(1)}(\pi_{1}(\Sigma))=H^{1}(\Sigma)$ by the Hurewicz theorem.
The construction of Frohman-Nicas has been subsequently extended to more general settings. In the work [6] by Bigelow, Cattabriga and Florens, the TQFT was extended to the category of tangles by quantizing the Burau representation of braids (see also [41, 42]). Complementarely, Florens and Massuyeau extended in [22] the Frohman-Nicas TQFT to the category of bordisms $W$ equipped with a representation $\pi_{1}(W)\to A$, for $A$ a fixed abelian group. Both constructions are essentially cohomological, being a a key tool the so-called Lescop’s Alexander function [39], which operates by considering exterior products of the Alexander module of a knot.
The TQFT developed by Frohman-Nicas and its extensions are inherently related to the ones constructed in Theorems B and C. Both constructions exploit the representation theory of the manifolds involved to quantize the Alexander polynomial, making use of the restriction maps onto the boundary. However, they are also very different in essence for several reasons. First, the Frohman-Nicas TQFT only uses abelian representations, whereas the non-abelianity of the group ${\textup{AGL}_{1}}(\mathbb{C})$ plays a fundamental role in our construction to capture the additive version of the Wirtinger relations. But, perhaps more importantly, the Frohman-Nicas construction is purely homological in nature, exploiting the homology theory of the involved manifolds, and thus linking with the Alexander polynomial through the intersection form. In sharp contrast, our TQFT is built upon subtle geometric properties of the ${\textup{AGL}_{1}}(\mathbb{C})$-representation theory, capturing the Alexander polynomial as the singular locus of a singular vector bundle. This also implies that our TQFT can be effectively computed from purely combinatorial information, which can be used to perform explicit calculations in highly involved knots.
It is an interesting prospective work to develop a general framework able to encompass both constructions, for instance, based on pull-push quantization on the derived category of coherent sheaves, à la Fourier-Mukai, as done for instance in [25, 57]. This potential relation can also be envisaged through the work of Cimasoni and Turaev [14, 15], where the authors use this homological machinery to construct another TQFT from the category of tangles into a category of “Lagrangian relations”, which might be seen as the homological analogue of our category of spans of singular vector bundles. In fact, the construction has been recently categorified in [13], where the authors construct a 2-functor from the category of tangles to the category of Lagrangian cospans. In contrast to the representation-theoretic framework explored in this manuscript, these approaches are homological in nature and do not involve representation varieties; computations have so far been carried out primarily under certain restrictions, such as topologically trivial tangles.
Finally, we want to emphasize that the quantization based on functorial field theory, as the one developed in the aforementioned works, differ from the ‘categorification’ approaches, such as Khovanov-like methods [36, 50], which capture polynomial invariants in a distinct way, namely, by constructing a homology theory whose (graded) Euler characteristic recovers the polynomial.
@@ 2. $\textup{AGL}_{1}$}-representations of knots
$\textup{(I)}\quad x_{i+1}=x_{j}x_{i}x_{j}^{-1},\qquad\textup{or}\qquad\textup{(II)}\quad x_{i+1}=x_{j}^{-1}x_{i}x_{j}\,$, (3)
@@
For simplicity, we will abbreviate ${\textup{AGL}_{1}}={\textup{AGL}_{1}}(\mathbb{C})$. Notice that algebraically, $\textup{AGL}_{1}$$\mathbb{C}$$^{*}\ltimes{\mathbb{C}}$, for the action of $\mathbb{C}$$^{*}$ on $\mathbb{C}$} by $t\cdot a=ta$.
@@
Analogously, each relation of type (II), $A_{j}^{-1}A_{i}A_{j}=A_{i+1}$, yields an analogous equation $a_{j}t^{-1}+a_{j}+t^{-1}a_{i}=a_{i+1}$, equivalent to
@@
Regarding non-abelian affine representations of ${\Gamma}_{K}$, the following fundamental theorem originally proven by de Rham and Burde in [17, Section 6.1] plays a fundamental role. We restate it in terms of the fibration given in (9), and we will re-prove it in Section 3.1.
@@ 3.1. Fox calculus for the Alexander module
Throughout this section, let us fix a knot $K\subset S^{3}$ with complement $K^{c}=S^{3}-K$. The commutator subgroup $\pi_{1}(K^{c}),\pi_{1}(K^{c})]\triangleleft\pi_{1}(K^{c})$ defines an infinite cyclic covering $K^{c}_{\infty}\rightarrow K^{c}$, with deck group isomorphic to $H_{1}(K^{c})=\pi_{1}(K^{c})/[\pi_{1}(K^{c}),\pi_{1}(K^{c})]\cong\mathbb{Z}=\langle t\rangle$, which is usually called the infinite cyclic covering of $K^{c}$. The cyclic action of $t$ endows each homology group $H_{k}(K^{c}_{\infty})$ with a $\mathbb{Z}[t,t^{-1}]$-module structure, giving rise to a collection of modules that are invariants of the knot, the Alexander modules of $K$.
@@
[…]
@@
The $k$-th Alexander polynomial $\Delta_{K}^{k}(t)$ of the knot $K$ is defined as the greatest common divisor of the elements of the ideal $E_{K}^{k}$, defined up to units $\pm t^{k}$, since ${\mathbb{Z}}[t,t^{-1}]$ is a unique factorization ring. In other words, $\Delta_{K}^{k}(t)$ is the generator of the smallest principal ideal containing $E_{K}^{k}$. Clearly, the first elementary ideal $E_{K}^{1}$ is a principal ideal, generated by a single polynomial $\Delta_{K}^{1}(t)$, also denoted by $\Delta_{K}(t)$, the Alexander polynomial, that appeared in Theorem 2.2.
@@
[35, Chapter 11] that $H_{1}(K^{c}_{\infty})$ is the ${\mathbb{Z}}[t,t^{-1}]$-module generated by $x_{i}],i=1,\ldots,n$, under the abelianized Wirtinger relations:
@@
according to the Wirtinger presentation of the knot. As it is also stated in [35, section 11.2], the linear relations (13), (14) in ${\mathbb{Z}}[t,t^{-1}]$ of the Alexander module are precisely the ones arising in any affine representation, from (7) and (8).
@@ 3.2. $\textup{AGL}_{1}$}-representation varieties and the Alexander module
${\mathcal{R}}_{{\textup{AGL}_{1}}}(K)]={\mathbb{L}}\left({\mathbb{L}}-1-\sum_{k=1}^{n-1}\left|S_{K}^{k}\right|\right)+\sum_{k=1}^{n-1}\left|S_{K}^{k}\right|{\mathbb{L}}^{k+1}={\mathbb{L}}({\mathbb{L}}-1)+\sum_{k=1}^{n-1}\left|S_{K}^{k}\right|({\mathbb{L}}^{k+1}-{\mathbb{L}}).$
@@ 3.3. The Alexander fibration as a singular vector bundle
${E_{\alpha}:=\pi^{-1}(U_{\alpha})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{\alpha}}$${U_{\alpha}\times\mathbb{C}^{n}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${U_{\alpha}}$
and such that the changes of charts are fibrewise linear. A morphism between singular vector bundles $\pi:E\to X$ and $\pi^{\prime}:E^{\prime}\to X^{\prime}$ is a pair (f,f~)($\tilde{f}$) of regular morphisms $f:X\to X^{\prime}$ and $\tilde{f}:E\to E^{\prime}$ such that the following diagram commutes
@@
Furthermore, as shown in [27, Proposition II.5.1], every coherent sheaf is uniquely determined by the finitely generated module $M_{K}=\mathcal{M}_{K}({\mathbb{C}}^{*})$ of global sections over the ring $\mathcal{O}_{{\mathbb{C}}^{*}}({\mathbb{C}}^{*})={\mathbb{C}}[t,t^{-1}]$. In fact, since $M_{K}$ is isomorphic to the ring of regular functions on ${\mathcal{R}}_{{\textup{AGL}_{1}}}(K)$, by Theorem 3.3 we get the following result.
@@
The coherent sheaf $\mathcal{M}_{K}$ has been previously studied by Hironaka in [30] in a different context, with an ad hoc construction by means of 11-cocycles. In this direction, the Alexander fibration (15) coincides with the one studied in [30].